OpenStudy (anonymous):

find the equation of a line with slope -6 and passing through (-2,-3)

myininaya (myininaya):

equation of a line has form: y=mx+b what does m represent?

OpenStudy (anonymous):

Use \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

Here m=-6 y1=-3 and x1=-2 Now substitute

OpenStudy (anonymous):

y+3=-6(x+2)

OpenStudy (anonymous):

how do you put it in ax+by=c

OpenStudy (anonymous):

now simplify y+3=-6x-12 6x+y=-15